OpenStudy (anonymous):

how many atoms of hydrogen are present in 25g of NH4Cl

OpenStudy (anonymous):

1 mole of NH4Cl = 14+4+35.5= 53.5 g 1 mole of NH4Cl = 4 moles of H 53.5 g of NH4Cl = 4 g of H 25 g of NH4Cl = 4/53.5 X 25 = 1.86 g of H